OpenStudy (anonymous):

Studying for the CLEP College Mathematics test and I found this question. I know the answer from a key, but I'm trying to understand how to solve it. Let 2^x = 16^(x-1). Which of the following is a solution? a) 1 + i and 1 - i b) 2 c) 1/2 + 3i and 1/2 - 3i d) 1 and -1

OpenStudy (anonymous):

I've been trying to solve by putting the solutions in for x, but I'm having some trouble dealing with the imaginary numbers. For instance, how to I solve 2^(1+i) = 16((1+i)-1)

OpenStudy (anonymous):

If someone could just talk me through solving it I would really appreciate it. I am new to open study but I'll help you however I can!

OpenStudy (kropot72):

\[\large 2^{x}=16^{x-1}\] \[\large 2^{x}=\frac{16^{x}}{16}\] \[\large 16=\frac{16^{x}}{2^{x}}=8^{x}\] \[\large x \log8=\log16\] \[\large x=\frac{\log16}{\log8}\] \[\large x=\frac{4}{3}\]

OpenStudy (anonymous):

The answer is actually b, which is 2. I understand what you did, but it's not correct?

Directrix (directrix):

The answer to the posted question is 4/3. I am thinking that the question or the options are both are incorrectly posted. To clear up the matter, would you post a screenshot of the actual question as it appears in your study materials. Thanks. @austenheroine29

OpenStudy (anonymous):

Sure -

Directrix (directrix):

The person who wrote the official solution made an error. See attachment and then ask questions about it if you like.

Directrix (directrix):